# Geometric Neural Network based on Phase Space for BCI-EEG decoding

### v2

@@ 5.4 Computational Performances
figurePerformance (ROC AUC) versus $CO_{2}$ consumption (g $CO_{2}$ equivalent). The size of the blob indicates the number of trainable parameters.
Avg. Time (s) $CO_{2}$ Emission (g $CO_{2}$)
DeepNet 21.77 0.39
ShallowNet 15.10 0.25
EEGNet-8,2 14.96 0.25
EEGITNet 28.44 0.48
EEGTCNet 28.83 0.47
EEGNeX 28.15 0.46
DynSpat+ShallowNet 34.42 0.41
DynSpat+EEGNet 34.09 0.45
SPDNet 6.80 0.07
Phase-SPDNet 30.16 0.22
@@
[…]
@@ 5.5 Limitation and Future Directions
The existing Phase-SPDNet , as employed thus far, utilizes MDOP for hyper-parameter selection, offering computational efficiency compared to grid-search methods [7]. However, the solutions obtained with MDOP were sub-optimal compared with using a more extensive hyper-parameter selection. A potential avenue for future exploration involves streamlining and accelerating the grid-search process to harness the capabilities of the Phase-SPDNet architecture fully.
@@
An intriguing direction arises from the insights gained in the explainability study, revealing that Phase-SPDNet predominantly focuses on diagonal elements. This suggests the possibility of employing a Region of Interest approach, where the network initially prioritizes the main diagonals while treating the remainder of the matrix as background.
@@
Another challenge was the difficulty in classifying data from less responsive subjects. Despite our efforts, there remains a need for improved methods better to capture the subjects’ intentions and their specific characteristics.
@@ 6 Conclusion
This study focuses on integrating the Augmented Covariance Matrix (ACM) with SPDNet to develop a practical algorithm intended for real-world applications, with a deliberate focus on scenarios involving a limited number of electrodes. Specifically, we restricted our electrode usage to just 3, strategically placed on the motor cortex.
In fact, the augmented methodology, based on Takens theorem, demonstrates a particularly effective performance when applied with a reduced number of electrodes. However, it is important to note that using only three electrodes can impact the results. For instance, in the BNCI2014001 dataset, employing just three electrodes led to a performance reduction of about $10\%$ in the ShallowNet architecture compared to using all available electrodes.
Validation of our approach involves leveraging nearly 100 subjects from several open datasets, a task facilitated by the MOABB framework. The resulting Phase-SPDNet algorithm outperforms existing DL algorithms in BCI-EEG classification and offers some explainability elements through GradCam++ visualization. Remarkably, this algorithm requires a modest number of trainable parameters and exhibits a lower environmental impact, measured by $CO_{2}$ equivalent emissions.
@@ 8 Open-source availability
The code will be made public during publication time under the BSD-3 License.
@@ Appendix A Results using Coherence as feature
In this section, we shift our focus to reporting results due to the use of Imaginary and Instantaneous coherence features as input. Specifically, we aim to draw a comparison between the Phase-SPDNet results and the non DL state-of-the-art pipelines (usually created for covariance but here applied on coherence). It’s worth noting that, in the subsequent analysis, we abstain from comparing these results with the performance of state-of-the-art DL models, given that the employment of coherence-based features typically yields lower results across the board. Nevertheless, we find it particularly intriguing to present the outcomes achieved by Phase-SPDNet in the context of coherence-based features since we observe that Phase-SPDNet demonstrates a statistically significant improvement. This shows the potential and significance of Phase-SPDNet in the broader landscape of signal processing and pattern recognition.
@@
Table A1: Summary of performances via average AUC-ROC (%) on six datasets (BCNI2014-01, BCNI2014-04, Cho2017, Schirrmeister2017, Weibo2014, and Zhou2016), for a left vs. right motor imagery task. Bold numbers represent the best score in each dataset. Estimators Imaginary Coherence and Instantaneous Coherence
@@
Models Estimator BNCI2014001 BNCI2014004 Cho2017 Schirrmeister2017 Weibo2014 Zhou2016 SPDNet Imn 56.55$\pm$7.41 58.12$\pm$12.20 56.16$\pm$8.94 59.15$\pm$7.84 60.59$\pm$10.78 62.89$\pm$8.97 CSP+LDA Imn 57.66$\pm$7.19 53.09$\pm$8.55 51.63$\pm$6.30 53.06$\pm$5.20 56.77$\pm$7.62 64.47$\pm$7.47 TANG+SVM Imn 59.39$\pm$8.86 53.66$\pm$8.26 52.45$\pm$6.24 52.93$\pm$4.83 56.09$\pm$8.89 64.19$\pm$9.49 FgMDM Imn 60.19$\pm$7.72 52.83$\pm$8.90 51.93$\pm$6.39 53.99$\pm$4.57 57.50$\pm$7.92 65.52$\pm$8.72 MDM Imn 60.30$\pm$7.83 54.04$\pm$9.00 52.09$\pm$6.67 53.36$\pm$5.34 57.31$\pm$7.33 66.27$\pm$8.60 Cov+EN Imn 60.71$\pm$7.92 52.67$\pm$8.96 52.07$\pm$6.70 54.42$\pm$4.07 57.21$\pm$8.23 64.58$\pm$8.41 Phase-SPDNet Imn 60.89$\pm$10.17 65.70$\pm$14.06 56.51$\pm$8.60 61.69$\pm$9.81 64.75$\pm$11.33 72.16$\pm$7.40 SPDNet Ins 63.42$\pm$5.76 63.71$\pm$12.88 57.61$\pm$11.09 59.73$\pm$6.89 66.11$\pm$12.02 75.07$\pm$7.63 MDM Ins 69.74$\pm$14.19 59.89$\pm$10.72 58.76$\pm$9.65 52.01$\pm$5.77 63.24$\pm$14.25 81.94$\pm$8.91 TANG+SVM Ins 69.82$\pm$14.76 60.44$\pm$9.76 58.86$\pm$10.11 53.31$\pm$5.42 63.10$\pm$15.51 85.02$\pm$8.29 CSP+LDA Ins 70.49$\pm$14.80 60.13$\pm$10.23 59.23$\pm$10.40 51.94$\pm$6.06 64.32$\pm$15.92 84.95$\pm$8.94 Cov+EN Ins 71.01$\pm$14.33 59.34$\pm$10.89 59.37$\pm$9.86 52.06$\pm$5.39 64.67$\pm$14.79 84.53$\pm$8.81 FgMDM Ins 71.61$\pm$13.86 60.64$\pm$10.24 59.45$\pm$10.29 52.53$\pm$5.32 65.09$\pm$14.73 85.17$\pm$8.36 Phase-SPDNet Ins 71.84$\pm$16.75 68.83$\pm$17.41 62.04$\pm$12.02 65.37$\pm$12.91 68.31$\pm$14.17 87.03$\pm$6.13
@@
In this study, the situation is even rosier: Phase-SPDNet , used with either imaginary or instantaneous coherence features, emerges as the best pipeline across all datasets. When utilizing imaginary coherence, SPDNet alone already achieves statistically superior results compared to state-of-the-art approaches employing the same features. However, the augmentation procedure introduces a significant performance boost, a finding supported by the meta-effect analysis shown in Fig A1 (c).
@@
In the case of instantaneous coherence, on the other hand, the impact of the augmentation procedure is even more significant. In fact, with this feature, the standard SPDNet fails to exhibit statistically superior performance in comparison to the state-of-the-art pipelines. Phase-SPDNet , on the other hand, stands out as the top-performing pipeline, showing that the augmentation procedure brings a statistically significant influence on classification outcomes (see Fig A2 (a)).
@@ Appendix B Additional Results
Figure A1: Result for right hand vs left hand classification, using within-session evaluation for state-of-the-art imaginary coherence pipelines. Plot (a) provides the relative improvement of the method considered with respect to the standard SPDNet of the different pipelines considered. Plot (b) shows a combined meta-analysis (over all datasets) of the different pipelines. It shows the significance of the algorithm on the y-axis being better than the one on the x-axis. The gray level represents the significance level of the ROC-AUC difference in terms of t-values. We only show significant interactions ($p<0.05$). Plots (c), (d), and (e) show the meta-analysis of Phase-SPDNet against SPDNet, COV+EN, and FgMDM, respectively. We show the standardized mean differences of p-values computed as a one-tailed Wilcoxon signed-rank test for the hypothesis given in the plot title. The gray bar denotes the $95\%$ interval. * stands for $p<0.05$, ** for $p<0.01$, and *** for $p<0.001$.
Figure A2: Result for right hand vs left hand classification, using within-session evaluation for state-of-the-art instantaneous coherence pipelines. Plot (a) provides the relative improvement of the method considered with respect to the standard SPDNet of the different pipelines considered. Plot (b) shows a combined meta-analysis (over all datasets) of the different pipelines. It shows the significance of the algorithm on the y-axis being better than the one on the x-axis. The gray level represents the significance level of the ROC-AUC difference in terms of t-values. We only show significant interactions ($p<0.05$). Plots (c), (d), and (e) show the meta-analysis of Phase-SPDNet against SPDNet, COV+EN, and FgMDM, respectively. We show the standardized mean differences of p-values computed as a one-tailed Wilcoxon signed-rank test for the hypothesis given in the plot title. The gray bar denotes the $95\%$ interval. * stands for $p<0.05$, ** for $p<0.01$, and *** for $p<0.001$.

### v3

@@ 5.4 Trainable Parameters
figurePerformance (ROC AUC) versus number of trainable parameters (logarithmic scale) for the Cho2017 dataset.
Stage Layer Output Parameters
Input [1, 3, T] 0
Augmentation AugmentedDataset [1, $3\times\psi$, $T-\psi\tau$] 0
Covariance Covariances [1, $\beta$, $\beta$] 0
SPDNet BiMap [1, $\alpha$, $\alpha$] $\alpha*\beta$
ReEig [1, $\alpha$, $\alpha$] 0
LogEig [1, $(\alpha*(\alpha+1))/2$] 0
Classification Linear [1, 2] $(\alpha*(\alpha+1))/2$
Total $(\alpha*\beta)+(\alpha*(\alpha+1))/2$
@@
tableArchitecture of Phase-SPDNet with parameter number depending on $\psi$.
Using $\beta=3\times\psi$ and $\alpha=(3\times\psi)/2$.
@@ 5.5 Limitation and Future Directions
An intriguing direction arises from the insights gained in the explainability study, revealing that Phase-SPDNet OPT predominantly focuses on diagonal elements. This suggests the possibility of employing a Region of Interest approach, where the network initially prioritizes the main diagonals while treating the remainder of the matrix as background.
@@
Another challenge was the difficulty in classifying data from less responsive subjects. Despite our efforts, there remains a need for improved methods to better capture the subjects’ intentions and their specific characteristics.
@@ 6 Conclusion
This study focuses on integrating the Augmented Covariance Matrix (ACM) with SPDNet to develop a practical algorithm intended for real-world applications, with a deliberate focus on scenarios involving a limited number of electrodes. Specifically, we restricted our electrode usage to just three, strategically placed on the motor cortex.
In fact, the ACM, based on Takens’ theorem, demonstrates a particularly effective performance when applied with a reduced number of electrodes. However, it is important to note that using only three electrodes can impact the results. For instance, in the BNCI2014001 dataset, employing just three electrodes led to a performance reduction of about $10\%$ in the ShallowNet architecture compared to using all available electrodes. We consider ShallowNet as it was the best performing DL model in this task [94].
Validation of our approach involves leveraging nearly 100 subjects from several open datasets, a task facilitated by the MOABB framework. The resulting Phase-SPDNet algorithm outperforms existing DL algorithms in BCI-EEG classification and offers some explainability elements through GradCam++ visualization. Remarkably, this algorithm requires a modest number of trainable parameters.
@@ 8 Open-source availability
The codes used to produce the results of this study are publicly available in this Github repository: https://github.com/carraraig/Phase-SPDNet.
@@ Appendix A Results using Coherence as feature
In this section, we report results using Imaginary and Instantaneous coherence features as input. Specifically, we aim to draw a comparison between the Phase-SPDNet , showing both MDOP and optimized hyperparameter search, results and the non DL state-of-the-art pipelines (usually created for covariance but here applied on coherence). It’s worth noting that, in the subsequent analysis, we abstain from comparing these results with the performance of state-of-the-art DL models, given that the employment of coherence-based features typically yields lower results across the board. Nevertheless, we find it particularly intriguing that both Phase-SPDNet OPT and Phase-SPDNet MDOP demonstrate a statistically significant improvement. This shows the potential and significance of Phase-SPDNet in the broader landscape of signal processing and pattern recognition. We outline below the pipeline used for comparison with the coherence features-based approach:
@@
Table A1: Summary of performances via average AUC-ROC (%) on six datasets (BCNI2014-01, BCNI2014-04, Cho2017, Schirrmeister2017, Weibo2014, and Zhou2016), for a left vs. right motor imagery task. Bold numbers represent the best score in each dataset. Estimators Imaginary Coherence (Imag) and Instantaneous Coherence (Inst)
@@
Models Estimator BNCI2014001 BNCI2014004 Cho2017 Schirrmeister2017 Weibo2014 Zhou2016 SPDNet Imag 56.55$\pm$7.41 58.12$\pm$12.20 56.16$\pm$8.94 59.15$\pm$7.84 60.59$\pm$10.78 62.89$\pm$8.97 CSP+LDA Imag 57.66$\pm$7.19 53.09$\pm$8.55 51.63$\pm$6.30 53.06$\pm$5.20 56.77$\pm$7.62 64.47$\pm$7.47 TANG+SVM Imag 59.39$\pm$8.86 53.66$\pm$8.26 52.45$\pm$6.24 52.93$\pm$4.83 56.09$\pm$8.89 64.19$\pm$9.49 FgMDM Imag 60.19$\pm$7.72 52.83$\pm$8.90 51.93$\pm$6.39 53.99$\pm$4.57 57.50$\pm$7.92 65.52$\pm$8.72 MDM Imag 60.30$\pm$7.83 54.04$\pm$9.00 52.09$\pm$6.67 53.36$\pm$5.34 57.31$\pm$7.33 66.27$\pm$8.60 Cov+EN Imag 60.71$\pm$7.92 52.67$\pm$8.96 52.07$\pm$6.70 54.42$\pm$4.07 57.21$\pm$8.23 64.58$\pm$8.41 Phase-SPDNet MDOP (Our) Imag 60.89$\pm$10.17 65.70$\pm$14.06 56.51$\pm$8.60 61.69$\pm$9.81 64.75$\pm$11.33 72.16$\pm$7.40 Phase-SPDNet OPT (Our) Imag 67.78$\pm$16.44 76.25$\pm$15.91 60.79$\pm$11.57 69.54$\pm$15.44 71.53$\pm$18.80 88.76$\pm$5.57 SPDNet Inst 63.42$\pm$5.76 63.71$\pm$12.88 57.61$\pm$11.09 59.73$\pm$6.89 66.11$\pm$12.02 75.07$\pm$7.63 MDM Inst 69.74$\pm$14.19 59.89$\pm$10.72 58.76$\pm$9.65 52.01$\pm$5.77 63.24$\pm$14.25 81.94$\pm$8.91 TANG+SVM Inst 69.82$\pm$14.76 60.44$\pm$9.76 58.86$\pm$10.11 53.31$\pm$5.42 63.10$\pm$15.51 85.02$\pm$8.29 CSP+LDA Inst 70.49$\pm$14.80 60.13$\pm$10.23 59.23$\pm$10.40 51.94$\pm$6.06 64.32$\pm$15.92 84.95$\pm$8.94 Cov+EN Inst 71.01$\pm$14.33 59.34$\pm$10.89 59.37$\pm$9.86 52.06$\pm$5.39 64.67$\pm$14.79 84.53$\pm$8.81 FgMDM Inst 71.61$\pm$13.86 60.64$\pm$10.24 59.45$\pm$10.29 52.53$\pm$5.32 65.09$\pm$14.73 85.17$\pm$8.36 Phase-SPDNet MDOP (Our) Inst 71.84$\pm$16.75 68.83$\pm$17.41 62.04$\pm$12.02 65.37$\pm$12.91 68.31$\pm$14.17 87.03$\pm$6.13 Phase-SPDNet OPT (Our) Inst 73.08$\pm$17.23 75.95$\pm$17.88 64.13$\pm$12.83 71.15$\pm$15.94 77.01$\pm$17.98 91.14$\pm$5.30
In this study, the situation is even rosier: Phase-SPDNet OPT, used with either imaginary or instantaneous coherence features, emerges as the best pipeline across all datasets. Interestingly, this study, which focuses on right versus left-hand motor imagery classification, ranks the information content in a ranking where the Imagination-based algorithm outperforms the instantaneous-based one. This result contrasts with the state-of-the-art understanding, which suggests that Instantaneous coherence mitigates the effects of volume conduction [51].
@@
When using imaginary coherence, SPDNet alone already achieves statistically superior results compared to state-of-the-art approaches employing the same features. However, the augmentation procedure introduces a significant performance boost, a finding supported by the meta-effect analysis shown in Fig A1 (c).
@@
In the case of instantaneous coherence, on the other hand, the impact of the augmentation procedure is even more significant. In fact, with this feature, the standard SPDNet fails to exhibit statistically superior performance in comparison to the state-of-the-art pipelines. Phase-SPDNet OPT, on the other hand, stands out as the top-performing pipeline, showing that the augmentation procedure brings a statistically significant influence on classification outcomes (see Fig A2 (a)).
@@ Appendix B Additional Results
Figure A1: Result for right hand vs left hand classification, using within-session evaluation for state-of-the-art imaginary coherence pipelines. Plot (a) provides the relative improvement of the method considered with respect to the standard SPDNet of the different pipelines considered. Plot (b) shows a combined meta-analysis (over all datasets) of the different pipelines. It shows the significance of the algorithm on the y-axis being better than the one on the x-axis. The gray level represents the significance level of the ROC-AUC difference in terms of t-values. We only show significant interactions ($p<0.05$). Plots (c), (d), and (e) show the meta-analysis of Phase-SPDNet OPT against SPDNet, COV+EN, and FgMDM, respectively. We show the standardized mean differences of p-values computed as a one-tailed Wilcoxon signed-rank test for the hypothesis given in the plot title. The gray bar denotes the $95\%$ interval. * stands for $p<0.05$, ** for $p<0.01$, and *** for $p<0.001$.
Figure A2: Result for right hand vs left hand classification, using within-session evaluation for state-of-the-art instantaneous coherence pipelines. Plot (a) provides the relative improvement of the method considered with respect to the standard SPDNet of the different pipelines considered. Plot (b) shows a combined meta-analysis (over all datasets) of the different pipelines. It shows the significance of the algorithm on the y-axis being better than the one on the x-axis. The gray level represents the significance level of the ROC-AUC difference in terms of t-values. We only show significant interactions ($p<0.05$). Plots (c), (d), and (e) show the meta-analysis of Phase-SPDNet OPT against SPDNet, COV+EN, and FgMDM, respectively. We show the standardized mean differences of p-values computed as a one-tailed Wilcoxon signed-rank test for the hypothesis given in the plot title. The gray bar denotes the $95\%$ interval. * stands for $p<0.05$, ** for $p<0.01$, and *** for $p<0.001$.